OpenStudy (anonymous):

The value of y varies directly with x2, and y = 150 when x = 5. What is the value of y when x = 3?

OpenStudy (freckles):

y varies directly with u means y=ku your u is x^2

OpenStudy (freckles):

you can find the constant k by using the point (5,150)

OpenStudy (freckles):

let me know if you need more help on this one

OpenStudy (anonymous):

please help

OpenStudy (freckles):

on which part?

OpenStudy (freckles):

can you show me what you have so far?

OpenStudy (freckles):

so I already said y=kx^2 and you are saying you don't know how to replace x with 5 and y with 150?

OpenStudy (freckles):

you replace the y with 150 and x with 5 like this y=kx^2 150=k(5)^2 solve for the constant k